For even integers x greater than 2, let *x* be defined as the product of the integers from 2 to x, inclusive. For example, *6*= (2)(4)(6)= 48.
If *8*/*6*=*a*, what is the value of a?

Answer by richard1234(7193)

*8*/*6* = (2)(4)(6)(8)/(2)(4)(6) = 8. This is equal to *2* = 2*4, so a = 2.